# ATOM: Attention Mixer for Efficient Dataset Distillation

### v1

@@ 4.1 Experimental Setup
Competitive Methods. In this paper, we compare the empirical results of ATOM on three computer vision datasets: CIFAR10/100 and TinyImageNet. We evaluate ATOM against four corset selection approaches and thirteen distillation methods for training set synthesis. The corset selection methods include Random selection [47], Herding [6, 4], K-Center [55], and Forgetting [60]. We also compare our approach with state-of-the-art distillation methods, including Dataset Distillation [63] (DD), Flexible Dataset Distillation [5] (LD), Dataset Condensation [76] (DC), Dataset Condensation with Contrastive (DCC) [38], Dataset Condensation with Differentiable Siamese Augmentation [74] (DSA), Distribution Matching [75] (DM), Deep Generative Priors (GLaD), Aligning Features [61] (CAFE), VIG [41], Kernel Inducing Points [44, 45] (KIP), Matching Training Trajectories [7] (MTT), and Attention Matching [51] (DAM).
@@ 4.2 Comparison with State-of-the-art Methods
Performance Comparison. In this section, we present a comparative analysis of our method against coreset and dataset distillation approaches. ATOM consistently outperforms these studies, especially at smaller distillation ratios, as shown in Table 1. Since the goal of dataset distillation is to generate a more compact synthetic set, we emphasize our significant performance improvements at low IPCs. We achieve almost $4\%$ improvement over the previous attention matching framework [51], DataDAM when evaluated on CIFAR-100 at IPC1. Notably, our performance on CIFAR-100 at IPC50 is 50.2% – that is nearly 90% of the baseline accuracy at a mere 10% of the original dataset. These examples motivate the development of dataset distillation works as downstream models can achieve relatively competitive performance with their baselines at a fraction of the training costs. Our primary objective in this study is to investigate the impact of channel-wise attention within the feature-matching process. Compared to prior attention-based and feature-based methodologies, our findings underscore the significance of channel-wise attention and the ATOM module, as validated also in the ablation studies in Figure 4.
@@
Distillation Cost Analysis. In this section, we delve into an examination of the training costs required for the distillation process. Although the main goal of dataset distillation is to reduce training costs across different applications such as neural architecture search and continual learning, the distillation technique itself must be efficient, enabling smooth operation on consumer-grade hardware. Approaches such as DC, DSA, and MTT introduce additional computational overhead due to bi-level optimization and training an expert model. In contrast, our method, akin to DM and DAM, capitalizes on randomly initialized networks, obviating the need for training and thereby reducing the computational cost per step involved in the matching stage. As illustrated in Table 3 utilizing solely the channel-based $\texttt{ATOM}^{\dagger}$ decreases the computational burden of matching compared to the default ATOM configuration. This efficiency is crucial, as channel-wise attention offers a more effective distillation process while maintaining superior performance (refer to Figure 4).

### v2

@@ 4.1 Experimental Setup
Competitive Methods. In this paper, we compare the empirical results of ATOM on three computer vision datasets: CIFAR10/100 and TinyImageNet. We evaluate ATOM against four corset selection approaches and thirteen distillation methods for training set synthesis. The corset selection methods include Random selection [47], Herding [6, 4], K-Center [55], and Forgetting [60]. We also compare our approach with state-of-the-art distillation methods, including Dataset Distillation [63] (DD), Flexible Dataset Distillation [5] (LD), Dataset Condensation [76] (DC), Dataset Condensation with Contrastive (DCC) [38], Dataset Condensation with Differentiable Siamese Augmentation [74] (DSA), Distribution Matching [75] (DM), Deep Generative Priors (GLaD), Aligning Features [61] (CAFE), VIG [41], Kernel Inducing Points [44, 45] (KIP) and Attention Matching [51] (DAM).
@@ 4.2 Comparison with State-of-the-art Methods
Performance Comparison. In this section, we present a comparative analysis of our method against coreset and dataset distillation approaches focusing on distribution matching related methods. ATOM consistently outperforms these studies, especially at smaller distillation ratios, as shown in Table 1. Since the goal of dataset distillation is to generate a more compact synthetic set, we emphasize our significant performance improvements at low IPCs. We achieve almost $4\%$ improvement over the previous attention matching framework [51], DataDAM when evaluated on CIFAR-100 at IPC1. Notably, our performance on CIFAR-100 at IPC50 is 50.2% – that is nearly 90% of the baseline accuracy at a mere 10% of the original dataset. These examples motivate the development of dataset distillation works as downstream models can achieve relatively competitive performance with their baselines at a fraction of the training costs. Our primary objective in this study is to investigate the impact of channel-wise attention within the feature-matching process. Compared to prior attention-based and feature-based methodologies, our findings underscore the significance of channel-wise attention and the ATOM module, as validated also in the ablation studies in Figure 4.
@@
Distillation Cost Analysis. In this section, we delve into an examination of the training costs required for the distillation process. Although the main goal of dataset distillation is to reduce training costs across different applications such as neural architecture search and continual learning, the distillation technique itself must be efficient, enabling smooth operation on consumer-grade hardware. Approaches such as DC, DSA introduce additional computational overhead due to bi-level optimization. Methods such as MTT achieve higher performance but incur the cost of training expert models in addition to overhead in the distillation process (see Table 3), for equitable comparison, we focus on distribution matching strategies. Within this domain, our method, akin to DM and DAM, capitalizes on randomly initialized networks, obviating the need for training and thereby reducing the computational cost per step involved in the matching stage. As illustrated in Table 3 utilizing solely the channel-based $\texttt{ATOM}^{\dagger}$ decreases the computational burden of matching compared to the default ATOM configuration. This efficiency is crucial, as channel-wise attention offers a more effective distillation process while maintaining superior performance (refer to Figure 4).